1/3

Christian Morgenstern,  Christian Morgenstern,  Christian Morgenstern,  Christian Morgenstern,

Le loup-garou Der Werwolf The Werewolf The Banshee (An Approach)

übersetzt von R. Platteau  übersetzt von Alexander Gross  übersetzt von Max Knight

Un loup-garou, une certaine nuit, Laissa sa moitié et puis se rendit Auprès de la tombe d’un maître d’école Et lui dit, “ sur moi dis tout, c’est ton rôle

! ”

Ein Werwolf eines Nachts ent-

wich von Weib und Kind und sich

begab an eines Dorfschullehrers Grab und bat ihn: �Bitte, beuge

mich!“ A Werewolf, troubled by his name, Left wife and brood one night and ca-

me To a hidden graveyard to enlist The aid of a long-dead philologist.

One night, a banshee slunk away from mate and child, and in the

gloom went to a village teacher’s tomb, requesting him: “Inflect me,

pray.”

Un loup-garou, une certaine nuit, Laissa sa moitié et puis se rendit Auprès de la tombe d’un maître d’école Et lui dit, “ sur moi dis tout, c’est ton rôle

! ”

Der Dorfschulmeister stieg hin-

auf auf seines Blechschilds Mes-

singknauf und sprach zum Wolf, der seine

Pfoten geduldig kreuzte vor dem To-

ten:

“Oh sage, wake up, please don’t berate

me,” He howled sadly, “Just conjugate me.” The seer arose a bit unsteady Yawned twice, wheezed once, and then

was ready.

The village teacher climbed up

straight upon his grave stone with its pla-

te and to the apparition said who meekly knelt before the

dead:

“ - On peut à ton sujet puisque nous y som-

mes Se demander encor bien des choses en som-

me. Loup-garoù ? mais avant Il faut se dire : loup-garquand ?

�Der Werwolf“ – sprach der gu-

te Mann, �des Werwolfs, Genitiv sodann, dem Werwolf, Dativ, wie man’s

nennt, den Werwolf, - damit hat’s ein

End!“ “Well, ‘Werewolf’ is your plural past, While ‘Waswolf’ is singularly cast: There’s ‘Amwolf’ too, the present ten-

se, And ‘Iswolf,’ ‘Arewolf’ in this same

sense.”

“The banSHEE, in the subject’s

place; the banHERS, the possessive ca-

se. The banHER, next, is what they

call objective case–and that is all.”

…  …  …  …

2/3

Loup-garcomment ? et puis Bien évidemment loup-garqui ? ” Toutes ces questions semblèrent au loup bi-

en absconses, Car en fait ce qu’il voulait c’était les

réponses.

Dem Werwolf schmeichelten

die Fälle er rollte seine Augenbälle. �Indessen“, bat er, �füge doch zur Einzahl auch die Mehrzahl

noch!“ “I know that–I’m no mental cripple– The future form and participle Are what I crave,” the beast replied. The scholar paused–again he tried:

The banshee marveled at the ca-

ses and writhed with pleasure, ma-

king faces, but said: “You did not add, so far, the plural to the singular!”

“ - Point de réponses, dit le puits de savoir, Etre imaginaire, nul ne peut te voir ! ” A ces mots l’animal éclata tout en larmes. “ - Oh!, même les baisers de ma douce moi-

tié Der Dorfschullehrer aber mus-

ste gestehn, dass er von ihr nichts

wusste. Zwar Wölfe gab’s in großer

Schar, doch �Wer“ gab’s nur im Sinu-

lar.

“A ‘Will-be-wolf?’ It’s just too long: ‘Shall-be-wolf?’ ‘Has-been-wolf?’ Ut-

terly wrong! Such words are wounds beyond all

suture– I’m sorry, but you have no future.”

The teacher, though, admitted

then that this was not within his ken. “While bans are frequent”, he ad-

vised, “A she cannot be plurized.”

Ne pourrons plus me faire me sentir tout

entier. Rester un loup-garou est un sort qui

m’alarme. ” Mais comme le loup était bien poli Il remercia et puis repartit.

Der Wolf erhob sich

tränenblind – er hatte ja doch Weib und

Kind! Doch da er keine Gelehrter

eben, so schied er dankend und erge-

ben.

The Werewolf knew better–his sons

still slept At home, and homewards now he

crept, Happy, humble, without apology For such folly of philology.

The banshee, rising clammily, wailed: “What about my family?” Then, being not a learned crea-

ture, said humbly “Thanks” and left

the teacher.

…  …  …  …

3/3

Übersetzung des deutschen Gedichtes “Der Werwolf” von Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) ins Französische durch R. Platteau.

Arg-825-1655 (2013-02-13 21:56:38)

Tiu ĉi francigo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Verfasser dieses deutschen Gedichtes ist Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31).

Arg-825-1652 (2013-02-11 18:44:21)

Übersetzung des deutschen Gedichtes “Der Werwolf” von Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) ins Englische durch Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Übersetzung des deutschen Gedichtes “Der Werwolf” von Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) ins Englische durch Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_term= m x% 20k% C3% BChnel&search_ country= DE&st= suc e% 20nach% 20personen&target_url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .